For the following reaction, 6.17 grams of diphosphorus pentoxide are mixed with excess water. The reaction yields 6.61 grams of phosphoric acid.


P2O5 (S) + 3H2O (I) ----> 2H3PO4 (aq)


(1) What is the theoretical yield of phosphoric acid? ..........grams


(2) What is the percent yield for this reaction? .........%


with steps please

1
Expert's answer
2021-03-02T01:02:11-0500

1 mole of P2O5 gives 2 mole of H3PO4

( 6.17/283.8 ) = 0.217 moles of P2O5

theoretically 0.0434 moles of H3PO4 should be produced

theoretical yield = 0.00434×97.9 = 4.24g

But it is 6.61 g which is might due to error in meaning

So % theoretical yield is 100%